OpenStudy (anonymous):

Solve for x: 2 over 5 (x - 4) = 2x. - The whole fraction thing really confuses me, someone walk me through this?

OpenStudy (anonymous):

First solve 5(x-4). That is 5x-20.

OpenStudy (anonymous):

Next, make both sides of the equation have the same denominator. If you can do that, the denominators will just disappear. So, the denominator on the left is 5x-2, and on the right is 1. To make both sides have the same denominator, multiply the right side by (5x-20)/(5x-20).

OpenStudy (anonymous):

Which gives you (on the right side) 2x(5x-20)/(5x-20).

OpenStudy (anonymous):

Drop the denominators (b/c they're equal on both sides) and you're left with 2=2x(5x-20).

OpenStudy (anonymous):

In my 2nd comment, I meant 5x-20 when I said 5x-2.

OpenStudy (anonymous):

@Avihirschx uh, the only answer choices I have are: -8 negative 1 over 2 1 -1

OpenStudy (anonymous):

I might have messed up something. Let me do the problem over again. Give me a minute.

OpenStudy (anonymous):

No, I was right. Perhaps I am confused as the how the problem looks? Is it this:\[\frac{ 2 }{ 5(x-4) }=2x\]

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 }(x-4)=2x\]

OpenStudy (anonymous):

@Avihirschx

OpenStudy (anonymous):

OK, that explains that.

OpenStudy (anonymous):

This can be expressed as: 2(x-4)/5=2x. Do you understand why?

OpenStudy (anonymous):

2/5(x-4)=2x 2(x-4)=2x.5 2x-8=10x 8x=-8 x=-1

OpenStudy (anonymous):

Yes. That^. I was trying to explain each step :).